Grade 12th passAlgebra

Find the area of the following circles, given that:
Diameter = 49 m

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given, diameter of circle (d) = 49 m
We know that, radius (r) = d/2
= 49/2
= 24.5 m
Then,
Area of the circle =πr2
= 22/7 × (24.5)2
= 22/7 × 600.25
= 22 × 85.75
= 1886.5 m2

Thanks